Find the range values for f(x) = -5x when the domain values are 1, 2, 3, and 4.
Answer by Hawksfan(61)

f(x) = -5x
another way to write this is
y= -5x
The domain is all the x values that you plug in.
The range is the y values or f(x) values.
For this problem, the domain is {1,2,3,4}.
f(1) = -5(1) = -5
f(2) = -5(2) = -10
f(3) = -5(3) = -15
f(4) = -5(4) = -20
therefore, the range is {-5,-10,-15,-20}